OpenStudy (anonymous):

Change 4x – 2y = 9 into slope-intercept form.

OpenStudy (cathyangs):

you just have to get it into y=mx+b, do you know how to do this?

OpenStudy (agreene):

4x – 2y = 9 subtract 4x from both sides -2y=9-4x divide by -2 y=9/(-2)-4x/(-2) y=2x-9/2

OpenStudy (cathyangs):

this is y=2x-9/2